The net accounts receivable on the books of GJY corporation as of January 1,2013 are as follows:

Account receivable

$562,000

Less: Allowance for sales discounts

($12,000)

Allowance for doubtful accounts

($41,300)

Net account receivable

$508,700

During the year, the sales discount allowance is left unchanged with discounts recorded directly in the sales discounts account. Allowances are adjusted at year-end. Summarized transactions during 2013 are as follows::

a) Sales revenue was $1,240,000 of which 70% was on credit.

b) Customers paid off $985,000 of the outstanding account receivable. Of this reduction to the account receivable, $600,000 was paid in time to earn a 2% discount.  

c) Accounts of $15,300 were written off at year-end.

d) There were recoveries of previously written-off accounts in the amount of $2,800.

e) At year-end, analysis of account receivable indicates that an allowance of $7,900 is needed for sales discounts.

f) At year-end, the allowance for doubtful accounts is increased by 1% of gross credit sales.

Required:

1) Prepare journal entries for the above transactions.

2) Show how the net accounts receivable appear on the statement of financial position on December 31,2013.

3) Do you have any concerns about the level of the allowance for doubtful accounts? Explain.
Answer #1
1 Particulars Debit Credit
a CASH $   3,72,000
ACCOUNTS RECEIVABLE $   8,68,000
SALES $ 12,40,000
B CASH $   9,85,000
Allowance for sales discounts $      12,000
ACCOUNTS RECEIVABLE $   9,97,000
C Allowance for doubtful accounts $      15,300
ACCOUNTS RECEIVABLE $       15,300
D CASH $         2,800
BAD DEBTS RECOVERED $         2,800
E SALES DISCOUNT $         7,900
Allowance for sales discounts $         7,900
f BAD DEBTS $         8,680
Allowance for doubtful accounts $         8,680
2 Particulars Amount
Account receivable $   5,62,000
Add: credit sales $   8,68,000
Less: Cash received $ -9,85,000
Less: allowance for sales discount $     -12,000
$   4,33,000
Less: Allowance for sales discounts $       -7,900
Less: Allowance for doubtful accounts $     -34,680
Net account receivable $   3,90,420

3. the percentage of allowance for doubtful accounts id 9% of net account receivables which is very high. Generally , the percentage is 2.5 %. It shows that the company is having high risk of non recoveries from accounts receivable.

Show formula in excel for calculations:

1 Particulars Debit Credit
a CASH =1240000*0.3
ACCOUNTS RECEIVABLE =1240000*0.7
SALES =+C3+C4
B CASH 985000
Allowance for sales discounts =600000*2%
ACCOUNTS RECEIVABLE =+C7+C8
C Allowance for doubtful accounts 15300
ACCOUNTS RECEIVABLE =+C11
D CASH 2800
BAD DEBTS RECOVERED =+C14
E SALES DISCOUNT 7900
Allowance for sales discounts 7900
f BAD DEBTS =1240000*0.7*0.01
Allowance for doubtful accounts =1240000*0.7*0.01
2 Particulars Amount
Account receivable 562000
Add: credit sales =1240000*0.7
Less: Cash received -985000
Less: allowance for sales discount =-600000*2%
=SUM(C25:C28)
Less: Allowance for sales discounts -7900
Less: Allowance for doubtful accounts =-41300+15300-8680
Net account receivable =SUM(C29:C31)